How much charge flows from each terminal of a 12.0-V battery when it is connected to a 5.00-μF capacitor? 


1
Expert's answer
2022-03-06T15:17:07-0500

The charge is given as the product of capacitance and voltage:

"Q= CV = 5\times 10^{-6}F\cdot 12V = 60.0\times 10^{-6}C"

Answer. "60.0\times 10^{-6}C".